simplymariex:

Suppose C and D represent two different school populations where C > D and C and D must be greater than 0. Which of the following expressions is the largest? Explain why. Show all work necessary. A: (C + D)^2 B: 2(C + D) C: C^2 + D^2 D: C^2 − D^2

Zepdrix:

Hey Marie :D Remember how to expand a square binomial? Like this! \(\large\rm (C+D)^2\quad=\quad C^2+2CD+D^2\)

Zepdrix:

If we expand out option B we get, \(\large\rm 2(C+D)\quad=\quad 2C+2D\)

simplymariex:

Ooh so it's option B?

Zepdrix:

Oh I left you hanging sorry :OOO

Zepdrix:

So this is A after we expand it out, \(\large\rm C^2+2CD+D^2\) And this is C, \(\large\rm C^2+D^2\) Since our values C and D are positive, you can clearly see how option A is larger, right? Option C is the same, but it's missing the middle part.

Zepdrix:

It would be easier to leave A in the non-expanded form so we can compare it to option B. So A looks like this, \(\large\rm (C+D)^2\) and B looks like this, \(\large\rm 2(C+D)\) As a general rule, square is going to result in a much much larger number than multiplying by 2. So therefore option A is larger than option B.

Zepdrix:

And option D is subtraction... so that's going to be a much smaller number. If they want you to justify it using some math techniques... you can actually factor option D into conjugates, \(\large\rm C^2-D^2\quad=\quad (C+D)(C-D)\) Which is smaller than option A, \(\large\rm (C+D)^2\quad=\quad (C+D)(C+D)\)

Falconmaster:

hey trust this guy ^ he know what he is talking about

SLAYEDQUEEN:

hey sarah